OpenStudy (anonymous):

a is now 4 years older than b and 2 years older than c. If in 2 years a will be twice as old as b,then in two years what would be a's age multilpied by c's age

OpenStudy (anonymous):

Let's set up some equations using what we know: "A is now 4 years older than B": \[a=b+4\] "A is ... 2 years older than C": \[a=c+2\] "in 2 years a will be twice as old as b" \[a+2=2(b+2)\] Then we are asked: "In two years what what is A's age times C'a age?" \[(a+2)(c+2)=?\] Good so far?

OpenStudy (anonymous):

We need 'a' and 'c'. To to this though, we will need 'b' first' We can find 'b' using equations 1 and 3: \[b+4=2(b+2)-2\]\[b=2(b+2)-2+4\]\[b=2b+4-2+4\]\[b=2b-2\]\[-b=-2\]\[b=2\] Using this we can get 'a':\[a=b+4\]\[a=2+4\]\[a=6\] and 'c':\[a=c+2\]\[6=c+2\]\[c=4\] And now solve the question: \[(a+2)(c+2)=?\]\[(6+2)(4+2)=?\]\[=48\]